OpenStudy (anonymous):

What are all the exact t-values fro which tan t = -√3? Multiple Choice a. π/3 - nπ b. -π/3+nπ c. π/5-nπ d. -π/5+nπ

OpenStudy (anonymous):

pi - pi/3, and 2pi - pi/3 are the two places that will give you the tan value. tan is - in quadrant 2 and 4. pi/3 values give sin = Root3/2 and cos= (1/2) tan = root of 3.